OpenStudy (anonymous):

5xy=1 using dy/dx the answer should be -x/y i just need to know hp to get there

OpenStudy (anonymous):

\[5y+5xy'=0\] \[5xy'=-5y\] \[y'=\frac{ -5y }{ 5x }\] \[y'=-\frac{ y }{ x }\]

OpenStudy (anonymous):

 actually is -y/x